OpenStudy (dls):

Find the angle between the parabolas y^2=4ax and x^2=4by at their point of intersection except origin.

OpenStudy (dls):

\[\LARGE \tan \theta= | \frac{m_1-m_2}{1+m_1m_2}|\]

OpenStudy (dls):

\[\LARGE 2y y'=4a =>y' = \frac{2a}{y}\] that is m1

OpenStudy (dls):

\[\LARGE 2x=4by' => y' = \frac{x}{2b}\] that is m2

OpenStudy (dls):

now I have to substitute x and y by solving both the curves together to get the value of m1 m2 right? but im getting weird things while trying to do that..

OpenStudy (dls):

If I'm proceeding correctly

OpenStudy (dls):

@ganeshie8

ganeshie8 (ganeshie8):

looks good, keep going :)

OpenStudy (dls):

getting double roots on x..

ganeshie8 (ganeshie8):

yes one must be 0 eh ? which you wanto discard anyways..

OpenStudy (dls):

yes one is 0 but dunno abt other

ganeshie8 (ganeshie8):

y^2 = 4ax -----(1) x^2 = 4by------(2) from (2), y = x^2/4b plug this in (1) (x^2/4b)^2 = 4ax solve x

OpenStudy (dls):

i did it wait hold on im getting weird stuff :/

ganeshie8 (ganeshie8):

(x^2/4b)^2 = 4ax x^4/16b^2 = 4ax x(x^3/16b^2-4a) = 0 x= 0, x^3/16b^2-4a = 0 x=0, x^3 = 64ab^2 x=0, x= \(4a^{1/3}b^{2/3}\)

OpenStudy (dls):

that is the weird stuff I meant! :O

OpenStudy (dls):

nvm

ganeshie8 (ganeshie8):

its not, its just constants a, b..

ganeshie8 (ganeshie8):

find y, yes i think this going to turn even more uglier... have fun ! :)